# Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths

### v1

@@ 4 Multi-wavelength Sérsic profiles
The HectoMAP data set enables investigation of trends in the rest-frame structural parameters as a function of wavelength for galaxies segregated in redshift, stellar mass, and average stellar population age. Sect. 4.1 outlines the procedure for measuring trends with rest-frame wavelength and the derived dependence on $z$ and $M_{\star}$. Sect. 4.2 investigates the changes in the wavelength dependence of Sérsic structural parameters (i.e. $R_{e}$ and $n$) as a function of $D_{n}4000$.
@@ 4.1 Variations in galaxy size and Sérsic index with wavelength
In the redshift range $0.60.2<z<$ 0.6, the rest-frame wavelengths ($\lambda^{RF}$) are up to $\sim 38\%$ smaller than the observed wavelengths ($\lambda_{f}$) in the five HSC bands ($f$). At $z=0.6$, the photometric bands correspond to rest-frame wavelengths that are $\sim 25\%$ bluer $\lambda^{RF}$ than at $z=0.2$. Light profiles covering a range of $\lambda^{RF}$ enable examination of the distributions of stellar populations with different ages within a galaxy. Generally the morphological parameters depend on $\lambda^{RF}$. The range of wavelengths spanned by the five bands of HSC, $\leavevmode\nobreak\ (4,000-10,700)\leavevmode\nobreak$ (Sect. 3.1) along with precise redshift determinations from Hectospec spectra enable determination of morphological parameters as a function of rest-frame wavelength..
@@
To obtain sizes, we set $\mathcal{P}=R_{e,c}[$kpc]]]], where $R_{e,c}$ is the circularized half-light radius. This radius averages out the galaxy asphericity; it is defined as $R_{e,c}=R_{e}\sqrt{b/a}$, where $b/a$ is the axis ratio. For Sérsic index, $\mathcal{P}=n$.
In Eq. (3) we consider differences between two rest-frame parameters $\mathcal{P}^{RF}$ at wavelengths $\lambda^{RF}_{1}=3500\leavevmode\nobreak$ and $\lambda^{RF}_{2}=7000\leavevmode\nobreak$. This choice allows exploration of the full dynamic range of observed wavelengths spanned by the HSC photometry without extrapolating beyond the observed range.
@@
The variable $\lambda_{f}$ includes the central observed wavelength of the five HSC bands: $\lambda_{f}=4750,6225,7700,8875$, and $10100\leavevmode\nobreak$ for the $g$-, $r$-, $i$-, $z$-, and $y$-band, respectively. Using HectoMAP spectroscopic redshifts we evaluate Eq. (4) at $\lambda_{1,2}=(1+z)\lambda^{RF}_{1,2}$ to estimate $R_{e,c}^{RF}$ and $n^{RF}$ in the two required rest frame wavelengths.
We use Eq. (3) to determine variations in $R_{e,c}$ and $n$ with wavelength. We examine the data in four equally-spaced bins of redshift and a variable number of equally-spaced $\Delta\left[\log(M_{\star}/M_{\odot})\right]=0.25$ dex bins in stellar mass. The number of stellar mass bins decreases with increasing redshift because the lower $M_{\star}$ limit increases with redshift. As a result, the number of mass bins decreases from eight to two from $0.30.2<z<$ 0.3 to $0.60.5<z<$ 0.6. At $0.30.2<z<$ 0.3, the six bins of increasing mass contain 996, 2432, 3244, 2660, 1212, and 319 galaxies, respectively. At $0.60.5<z<$ 0.6, the two mass bins contain 1645 and 1245 galaxies. Overall, the mass bins contain a statistically large number of galaxies regardless of redshift and corresponding mass limit. The sample size is similar for the different bins.
@@
Circles in Figure 3 show the median rate of change in sizes $R_{e,c}$ with rest-frame wavelength. Size always decreases with increasing wavelength, in the range $\sim-(0.1-0.3)$; $R_{e,c}$ decreases with increasing $\lambda^{RF}$. In other words, as is well-known from other studies, the cores of quiescent galaxies are redder than their outer regions.
@@
The absolute value of the size changes with wavelength at a rate that depends mildly on redshift. The median $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}$ is $\sim-0.3$ for $z>0.4$, and $\sim-0.1$ for $0.30.2<z<$ 0.3, implying an average $\sim 30\%$ decrease in absolute value as the redshift decreases from $z\sim 0.4$ to $z\sim 0.2$. The interquartile range (shadowed areas in Figure 3) shows that the relative scatter is $\sim 100\%$ at any redshift. These changes in size are independent of $M_{\star}$.
@@
The HectoMAP-based estimates of $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}$ at $0.40.2<z<$ 0.4 are within the range of values based on statistical spectroscopic samples at $z\lesssim 0.25$ ($\sim-0.16$, GAMA survey, Kelvin et al. 2012; Kennedy et al. 2016) and much smaller spectroscopic samples from a broader redshift interval ($-0.25$ for $20<z<$ 2, 3D-HST+CANDELS, van der Wel et al. 2014). A recent photometric study exploring the redshift range of HectoMAP also reports similar results for $\sim 10^{11}\,M_{\sun}$ quiescent systems, with $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}\sim-(0.10-0.23)$ at $z\sim 0.3$, and $\sim-(0.29-0.40)$ at $z\sim 0.6$ (Table 3 from George et al. 2024).
Figure 4: Same as Figure 3, but for Sérsic index.
@@
The median rate of change in Sérsic index with stellar mass and redshift (Figure 4) shows that the light profiles of quiescent galaxies are always more centrally concentrated at longer rest-frame wavelengths (i.e., $\frac{\Delta\log n}{\Delta\log\lambda^{RF}}$ is always positive). The rate of change in $n^{RF}$ with rest-frame wavelength exhibits is essentially independent of redshift or $M_{\star}$. The average $\frac{\Delta\log n}{\Delta\log\lambda^{RF}}$ is $\sim 0.20$ for $0.50.2<z<$ 0.5, and $\sim 0.3$ for $0.60.5<z<$ 0.6. The relative dispersions are $\sim 100\%$ independent of redshift and $M_{\star}$.
@@
The sizes of quiescent galaxies at $0.60.2<z<$ 0.6 are larger at bluer wavelengths (Figure 3). In contrast, their median Sérsic index increases with wavelength (Figure 4). The red light profiles that trace the bulk of the galaxy stellar mass distribution are smaller and more centrally concentrated than the blue light profiles dominated by younger or more metal-poor stellar populations. The wavelength-dependent properties of light profiles in the HectoMAP sample confirm the results from spatially resolved photometric studies based either on statistically large samples of quiescent systems at $z\lesssim 0.2$ (e.g., Peletier et al. 1990; Tamura & Ohta 2000; La Barbera et al. 2005; La Barbera & de Carvalho 2009; Gonzalez-Perez et al. 2011) or significantly smaller samples at $z\lesssim 2$ (e.g., Guo et al. 2011; Gargiulo et al. 2012; Tacchella et al. 2015).
@@
The fact that the cores of quiescent systems are redder than their outer regions has the potential to constrain quenching within galaxies (i.e, inside-out quenching, where star formation terminates first in the central regions, Avila-Reese et al. 2018; Nelson et al. 2021; McDonough et al. 2023; Mun et al. 2024, 2025). However, galaxies grow in size during the quiescent phase and this growth is also wavelength-dependent (e.g., Naab et al. 2009; Hilz et al. 2012; Oser et al. 2012, Section 5.2). To distinguish between the effects of the two processes, it is critical to compare the wavelength dependence of structural parameters for galaxies with different average stellar population ages.
@@
Newcomers are galaxies that recently joined the quiescent population. We identify newcomers at each redshift as galaxies with $1.61.5<D_{n}4000<$ 1.6 (Vergani et al. 2008; Damjanov et al. 2018; Hamadouche et al. 2022; Damjanov et al. 2023). There are 3888 newcomers in the mass-complete sample. Aging galaxies have $D_{n}4000>1.5$ at $z\sim 0.6$. At lower redshift, aging galaxies have larger $D_{n}4000$ values because of the relation between this spectral index and the galaxy global stellar population age.
@@
In Damjanov et al. (2023) we develop a procedure for selecting the aging population in the HectoMAP sample based on $D_{n}4000$. Using the Flexible Stellar Population Synthesis code (Conroy et al. 2009; Conroy & Gunn 2010), we derive evolving synthetic spectra for galaxies that by $z=0.6$ have not been forming stars for 0, 1, 2, 3, or 4 Gyr. Redshift evolution from $z\sim 0.6$ to $z\sim 0.2$ of these synthetic spectra is based on the increase in $D_{n}4000$ as the universe ages over this redshift range. The general increase in $D_{n}4000$ with decreasing redshift enable selection of $0.60.2<z<$ 0.6 HectoMAP galaxies with average stellar population ages equivalent to the descendants of quiescent $z\sim 0.6$ galaxies. There are 13,564 aging quiescent galaxies in our mass-complete sample.
@@
Figure 5: $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}$ as a function of stellar mass and $D_{n}4000$ for two quiescent subpopulations. Circles show the median rate of change in size with rest-frame wavelength for the subpopulations of aging galaxies (shades of purple and red) and newcomers (blue) in $\Delta\left[\log(M_{\star}/M_{\odot})\right]=0.25$ bins of stellar mass. The color of the symbols corresponds to the median $D_{n}4000$ index of the subpopulation. The panels show the four $\Delta z=0.1$ intermediate redshift bins (as indicated in the panel legends). Error bars denote the interquartile range of the correspondingly colored population. Colors indicate $D_{n}4000$. The black curve in each panel shows the median rate of change in size with rest-frame wavelength of the total quiescent population (same as Figure 3) and the gray area shows the corresponding interquartile range.
@@
Figure 5 illustrates the dependence of variations in galaxy size with rest-frame wavelength on the median $D_{n}4000$ index that separates newcomers (blue circles) from the aging quiescent population (purple- and red-shaded circles) in four $\Delta z=0.1$ redshift bins with $0.60.2<z<$ 0.6. In each redshift bin, the subpopulations are separated into $\Delta\left[\log(M_{\star}/M_{\odot})\right]=0.25$ stellar mass bins.
@@
For all stellar mass bins covering $0.30.2<z<$ 0.3, $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}$ for newcomers is $\sim 10-50\%$ larger in absolute value than for the aging population. The difference between the two subpopulations increases steadily with increasing $M_{\star}$. There is no difference between the variations in size with rest-frame wavelength for newcomers and the aging population for $z>0.3$.
@@
At $0.30.2<z<$ 0.3 newcomers have a systematically smaller ratio between their sizes at redder and bluer wavelengths compared to aging galaxies. Relative to their stellar mass distribution, newcomers at this redshift have more extended blue light profiles than the aging subpopulation.
Figure 6: Same as Figure 5, but for $\frac{\Delta\log n}{\Delta\log\lambda^{RF}}$.
@@
Figure 6 shows the variations in Sérsic index with wavelength as a function of stellar mass for two quiescent populations. At fixed stellar mass, there is a clear decrease in $\frac{\Delta\log n}{\Delta\log\lambda^{RF}}$ with galaxy global stellar population age. The median rate of increase in Sérsic index with wavelength for newcomers is $\sim 3,4,2$, and 1.5 times larger than for the aging population at $z\sim 0.25,0.35,0.45,$ and $0.55$, respectively. In all but the highest mass bin at $z<0.4$ (first two panels of Figure 6), extremely low $p-$values of the KS tests ($\lesssim 10^{-7}$) confirm that this difference in the increase of Sérsic index with wavelength between two quiescent populations is statistically significant. [^3] Equivalent to the difference in the wavelength-dependence of size between newcomers and aging population, the observed change in the increase of Sérsic index with wavelength between the two subpopulations is not statistically significant at $z>0.4$ because the sample sizes and age differences are small.
@@
At $0.40.2<z<$ 0.4, galaxies that only recently ceased star formation exhibit a systematically larger ratio between the central concentrations of their red and UV light profiles with respect to the aging galaxies. Relative to the distribution of their stellar mass, newcomers are less centrally concentrated in UV light than the resident quiescent galaxies that ceased their star formation activity at earlier epochs.
@@
The changes in morphological parameters with the rest-frame wavelength depend on the stellar population age and thus provide additional constraints on quenching processes. The ratio between the Sérsic indices in the red and UV rest-frames exceeds unity for the newcomers and for the aging population. However, newcomers to the quiescent population show larger ratios than the resident (aging) population in all stellar mass and redshift bins (Figure 6). At $0.30.2<z<$ 0.3, where the difference in age between the old (aging) population and young quiescent galaxies (newcomers) is most significant, the ratio between sizes of galaxies in the red and the UV rest-frame is also systematically larger for the population of newcomers. In this redshift bin, light profiles (i.e., $R_{e,c}$ and $n$) at two rest-frame wavelengths - $3500$ and $7000$ are more similar for the aging population than for the newcomers.

### v2

@@ 4 Multiwavelength Sérsic profiles
The HectoMAP data set enables the investigation of trends in the rest-frame structural parameters as a function of wavelength for galaxies segregated in redshift, stellar mass, and average stellar population age. Section 4.1 outlines the procedure for measuring trends with rest-frame wavelength and the derived dependence on $z$ and $M_{\star}$. Section 4.2 investigates the changes in the wavelength dependence of Sérsic structural parameters (i.e. $R_{e}$ and $n$) as a function of $D_{n}4000$.
@@ 4.1 Variations in galaxy size and Sérsic index with wavelength
In the redshift range $0.2<z<0.6$, the rest-frame wavelengths ($\lambda^{RF}$) are up to $\sim 38\%$ smaller than the observed wavelengths ($\lambda_{f}$) in the five HSC bands ($f$). At $z=0.6$, the photometric bands correspond to rest-frame wavelengths that are $\sim 25\%$ bluer $\lambda^{RF}$ than at $z=0.2$. Light profiles covering a range of $\lambda^{RF}$ enable examinations of the distributions of stellar populations with different ages within a galaxy. Generally the morphological parameters depend on $\lambda^{RF}$. The range of wavelengths spanned by the five bands of HSC, $\penalty 10000\ (4,000-10,700)\penalty 10000\$Å (Sect. 3.1) along with precise redshift determinations from Hectospec spectra enable the determination of morphological parameters as a function of rest-frame wavelength.
@@
To obtain sizes, we set $\mathcal{P}=R_{e,c}[$kpc]], where $R_{e,c}$ is the circularized half-light radius. This radius averages out the galaxy asphericity; it is defined as $R_{e,c}=R_{e}\sqrt{b/a}$, where $b/a$ is the axis ratio. For Sérsic index, $\mathcal{P}=n$.
In Eq. (2), we consider the differences between two rest-frame parameters, $\mathcal{P}^{RF}$, at wavelengths of $\lambda^{RF}_{1}=3500\penalty 10000\$Å and $\lambda^{RF}_{2}=7000\penalty 10000\$Å. This choice enables the exploration of the full dynamic range of observed wavelengths spanned by the HSC photometry without extrapolating beyond the observed range.
@@
The variable $\lambda_{f}$ includes the central observed wavelength of the five HSC bands: $\lambda_{f}=4750,6225,7700,8875$, and $10100\penalty 10000\$Å for the $g$, $r$, $i$, $z$, and $y$ bands, respectively. Using HectoMAP spectroscopic redshifts we evaluate Eq. (3) at $\lambda_{1,2}=(1+z)\lambda^{RF}_{1,2}$ to estimate $R_{e,c}^{RF}$ and $n^{RF}$ in the two required rest frame wavelengths.
We use Eq. (2) to determine variations in $R_{e,c}$ and $n$ with wavelength. We examine the data in four equally spaced bins of redshift and a variable number of equally spaced $\Delta\left[\log(M_{\star}/M_{\odot})\right]=0.25$ dex bins in stellar mass. In Table 3 we report the total number of galaxies in our four redshift bins and the corresponding range of lower limits on $\log(M_{\star}/M_{\odot})$ for the mass-complete sample (blue line in the central panel of Fig. 1). The total number of galaxies decreases with increasing redshift. However, the sample size remains statistically large in all four redshift bins.
Table 3: Sampling and mass limit for the HectoMAP redshift bins.
$z$ range no. of galaxies $\log(M_{\ast,lim}/M_{\odot})$
$0.2<z<0.3$ 10863 (9.66-10.11)
$0.3<z<0.4$ 8439 (10.11-10.51)
$0.4<z<0.5$ 4720 (10.59-10.87)
$0.5<z<0.6$ 2890 (10.87-11.18)
Notes. The lower and upper limit of the intervals in $\log(M_{\star,lim}/M_{\odot})$ correspond to the values of the $M_{\star,lim}(z)$ function (blue curve in Fig. 1) at the lower and upper edges of the four redshift bins from the first column.
The lower and upper masses sampled in our mass bins are $\log(M_{\star},min/M_{\odot})=10.0$ and $\log(M_{\star},max/M_{\odot})=11.5$, respectively. This choice enables us to include a minimum of 300 galaxies in each equally spaced $\Delta\left[\log(M_{\star}/M_{\odot})\right]=0.25$ dex bin in stellar mass at any redshift. The number of stellar mass bins decreases with increasing redshift because the lower $M_{\star}$ limit increases with redshift. As a result, the number of mass bins decreases from eight to two from $0.2<z<0.3$ to $0.5<z<0.6$. At $0.2<z<0.3$, the six bins of increasing mass contain 996, 2432, 3244, 2660, 1212, and 319 galaxies, respectively. At $0.5<z<0.6$, the two mass bins contain 1645 and 1245 galaxies. Overall, the mass bins contain a statistically large number of galaxies regardless of redshift and corresponding mass limit.
@@
Circles in Fig. 3 show the median rate of change in sizes $R_{e,c}$ with rest-frame wavelength. Size always decreases with increasing wavelength, in the range $\sim-(0.1-0.3)$; $R_{e,c}$ decreases with increasing $\lambda^{RF}$. In other words, as is well-known from other studies, the cores of quiescent galaxies are redder than their outer regions.
@@
The absolute value of the size changes with wavelength at a rate that depends mildly on redshift. The median $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}$ is $\sim-0.3$ for $z>0.4$, and $\sim-0.1$ for $0.2<z<0.3$, implying an average $\sim 30\%$ decrease in absolute value as the redshift decreases from $z\sim 0.4$ to $z\sim 0.2$. The interquartile range (shadowed areas in Fig. 3) shows that the relative scatter is $\sim 100\%$ at any redshift. These changes in size are independent of $M_{\star}$.
@@
The HectoMAP-based estimates of $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}$ at $0.2<z<0.4$ are within the range of values based on statistical spectroscopic samples at $z\lesssim 0.25$ ($\sim-0.16$, GAMA survey, Kelvin et al. 2012; Kennedy et al. 2016) and much smaller spectroscopic samples from a broader redshift interval (0.25-0.25 for $0<z<2$, 3D-HST+CANDELS, van der Wel et al. 2014). A recent photometric study exploring the redshift range of HectoMAP also reports similar results for $\sim 10^{11}\,M_{\sun}$ quiescent systems, with $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}\sim-(0.10-0.23)$ at $z\sim 0.3$, and $\sim-(0.29-0.40)$ at $z\sim 0.6$ (Table 3 from George et al. 2024).
Figure 4: Same as Fig. 3 but for the Sérsic index.
@@
The median rate of change in Sérsic index with stellar mass and redshift (Fig. 4) shows that the light profiles of quiescent galaxies are always more centrally concentrated at longer rest-frame wavelengths (i.e., $\frac{\Delta\log n}{\Delta\log\lambda^{RF}}$ is always positive). The rate of change in $n^{RF}$ with rest-frame wavelength is essentially independent of redshift or $M_{\star}$. The average $\frac{\Delta\log n}{\Delta\log\lambda^{RF}}$ is $\sim 0.20$ for $0.2<z<0.5$, and $\sim 0.3$ for $0.5<z<0.6$. The relative dispersions are $\sim 100\%$ independent of redshift and $M_{\star}$.
@@
The sizes of quiescent galaxies at $0.2<z<0.6$ are larger at bluer wavelengths (Fig. 3). In contrast, their median Sérsic index increases with wavelength (Fig. 4). The red light profiles that trace the bulk of the galaxy stellar mass distribution are smaller and more centrally concentrated than the blue light profiles dominated by younger or more metal-poor stellar populations. The wavelength-dependent properties of light profiles in the HectoMAP sample confirm the results from spatially resolved photometric studies based either on statistically large samples of quiescent systems at $z\lesssim 0.2$ (e.g., Peletier et al. 1990; Tamura & Ohta 2000; La Barbera et al. 2005; La Barbera & de Carvalho 2009; Gonzalez-Perez et al. 2011) or significantly smaller samples at $z\lesssim 2$ (e.g., Guo et al. 2011; Gargiulo et al. 2012; Tacchella et al. 2015).
@@
The fact that the cores of quiescent systems are redder than their outer regions has the potential to constrain quenching within galaxies (i.e, inside-out quenching, where star formation terminates first in the central regions, Avila-Reese et al. 2018; Nelson et al. 2021; McDonough et al. 2023; Mun et al. 2024, 2025). However, galaxies grow in size during the quiescent phase and this growth is also wavelength-dependent (e.g., Naab et al. 2009; Hilz et al. 2012; Oser et al. 2012, Sect. 5.2). To distinguish between the effects of the two processes, it is critical to compare the wavelength dependence of structural parameters for galaxies with different average stellar population ages.
@@
Newcomers are galaxies that recently joined the quiescent population. We identify newcomers at each redshift as galaxies with $1.5<D_{n}4000<1.6$, a standard selection in spectroscopic studies at $z\lesssim 1$ (Kauffmann et al. 2003; Vergani et al. 2008; Damjanov et al. 2023; Figueira et al. 2024). There are 3888 newcomers in the mass-complete sample. Aging galaxies have $D_{n}4000>1.5$ at $z\sim 0.6$. At lower redshift, aging galaxies have larger $D_{n}4000$ values because of the relation between this spectral index and the galaxy global stellar population age.
@@
In Damjanov et al. (2023), we developed a procedure for selecting the aging population in the HectoMAP sample based on $D_{n}4000$. Using the Flexible Stellar Population Synthesis code (FSPS, Conroy et al. 2009; Conroy & Gunn 2010), we derived evolving synthetic spectra for galaxies that by $z=0.6$ have not been forming stars for 0, 1, 2, 3, or 4 Gyr. Redshift evolution from $z\sim 0.6$ to $z\sim 0.2$ of these synthetic spectra is based on the increase in $D_{n}4000$ as the universe ages over this redshift range. The general increase in $D_{n}4000$ with decreasing redshift enable the selection of $0.2<z<0.6$ HectoMAP galaxies with average stellar population ages equivalent to the descendants of quiescent $z\sim 0.6$ galaxies. There are 13,564 aging quiescent galaxies in our mass-complete sample.
@@
Figure 5: $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}$ as a function of stellar mass and $D_{n}4000$ for two quiescent subpopulations. Circles show the median rate of change in size with rest-frame wavelength for the subpopulations of aging galaxies (shades of purple and red) and newcomers (blue) in $\Delta\left[\log(M_{\star}/M_{\odot})\right]=0.25$ bins of stellar mass. The color of the symbols corresponds to the median $D_{n}4000$ index of the subpopulation. The panels show the four $\Delta z=0.1$ intermediate redshift bins (as indicated in the panel legends). Error bars denote the interquartile range of the correspondingly colored population. Colors indicate $D_{n}4000$. The black curve in each panel shows the median rate of change in size with rest-frame wavelength of the total quiescent population (same as Fig. 3) and the gray area shows the corresponding interquartile range.
@@
Figure 5 illustrates the dependence of variations in galaxy size with rest-frame wavelength on the median $D_{n}4000$ index that separates newcomers (blue circles) from the aging quiescent population (purple- and red-shaded circles) in four $\Delta z=0.1$ redshift bins with $0.2<z<0.6$. In each redshift bin, the subpopulations are separated into $\Delta\left[\log(M_{\star}/M_{\odot})\right]=0.25$ stellar mass bins.
@@
For all stellar mass bins covering $0.2<z<0.3$, $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}$ for newcomers is $\sim 10-50\%$ larger in absolute value than for the aging population. The difference between the two subpopulations increases steadily with increasing $M_{\star}$. There is no difference between the variations in size with rest-frame wavelength for newcomers and the aging population for $z>0.3$.
@@
At $0.2<z<0.3$ newcomers have a systematically smaller ratio between their sizes at redder and bluer wavelengths compared to aging galaxies. Relative to their stellar mass distribution, newcomers at this redshift have more extended blue light profiles than the aging subpopulation.
Figure 6: Same as Fig. 5 but for $\frac{\Delta\log n}{\Delta\log\lambda^{RF}}$.
@@
Figure 6 shows the variations in Sérsic index with wavelength as a function of stellar mass for two quiescent populations. At fixed stellar mass, there is a clear decrease in $\frac{\Delta\log n}{\Delta\log\lambda^{RF}}$ with galaxy global stellar population age. The median rate of increase in Sérsic index with wavelength for newcomers is $\sim 3,4,2$, and 1.5 times larger than for the aging population at $z\sim 0.25,0.35,0.45,$ and $0.55$, respectively. In all but the highest mass bin at $z<0.4$ (first two panels of Fig. 6), extremely low $p$-values of the KS tests ($\lesssim 10^{-7}$) confirm that this difference in the increase of Sérsic index with wavelength between two quiescent populations is statistically significant. [^3] Equivalent to the difference in the wavelength-dependence of size between newcomers and aging population, the observed change in the increase of Sérsic index with wavelength between the two subpopulations is not statistically significant at $z>0.4$ because the sample sizes and age differences are small.
@@
At $0.2<z<0.4$, galaxies that only recently ceased star formation exhibit a systematically larger ratio between the central concentrations of their red and UV light profiles with respect to the aging galaxies. Relative to the distribution of their stellar mass, newcomers are less centrally concentrated in UV light than the resident quiescent galaxies that ceased their star formation activity at earlier epochs.
@@
The changes in morphological parameters with the rest-frame wavelength depend on the stellar population age and thus provide additional constraints on quenching processes. The ratio between the Sérsic indices in the red and UV rest frames exceeds unity for the newcomers and for the aging population. However, newcomers to the quiescent population show larger ratios than the resident (aging) population in all stellar mass and redshift bins (Fig. 6). At $0.2<z<0.3$, where the difference in age between the old (aging) population and young quiescent galaxies (newcomers) is most significant, the ratio between sizes of galaxies in the red and the UV rest-frame is also systematically larger for the population of newcomers. In this redshift bin, light profiles (i.e., $R_{e,c}$ and $n$) at two rest-frame wavelengths - $3500$ and $7000$ Å - are more similar for the aging population than for the newcomers.
We explore the effect of $D_{n}4000$ measurement errors ($\Delta(D_{n}4000)=0.067$ on average; Sect. 2.1) on the differences in $R_{e}$ and $n$ variations with wavelength between newcomers and aging galaxies by repeating the analysis using 1.67 (instead of 1.6) as the upper limit of the $D_{n}4000$ interval for newcomers. In stellar mass and redshift bins of Figs. 5 and 6, the rates of decrease in $R_{e}$ and increase in $n$ with wavelength for newcomers differ on average by 1.5% and 4.9%, respectively. These differences are well within the statistical uncertainties of our $\frac{\Delta\log R_{e,c}}{\Delta\log\lambda^{RF}}$ and $\frac{\Delta\log n}{\Delta\log\lambda^{RF}}$estimates. The variations in structural parameters with wavelength for the aging population are not affected.